OpenStudy (anonymous):

write 55% as a fraction

OpenStudy (anonymous):

55/100 then reduce

OpenStudy (anonymous):

11/20

OpenStudy (anonymous):

thats the reduced form

OpenStudy (mathstudent55):

Percent means "out of 100" To write a percent as a fraction, just divide by 100. 55% = 55/100 = 11/20

OpenStudy (anonymous):

so its 11/55

OpenStudy (mathstudent55):

11/20 is the answer

OpenStudy (anonymous):

If you do 11/20 in your calculator you will get .55 which is the decimal form for 55%

OpenStudy (anonymous):

its not the answer kuz its not an option

OpenStudy (mathstudent55):

Is 55/100 an option?